OpenStudy (anonymous):

Find the two products below. Compare and contrast, in complete sentences, the similarities and differences of the two. (x + 4)(x − 4) and (x + 4)(x + 4)

OpenStudy (anonymous):

1) x^2-16 2)x^2+8x+16 The first product cancelled the +8x, and instead of making the 16 positive, it made it negative. this is because of the face that when expanded (x+4)(x-4)= x^2+4x-4x-16 while (x+4)(x+4) = x^2+4x+4x+16.

OpenStudy (anonymous):

So, the similarities and differences are non exisitent?

OpenStudy (anonymous):

do you mean the similarities in the graphs?

OpenStudy (anonymous):

It asks for the similarities and differences

OpenStudy (anonymous):

Similarity: they are both parabolas Differences: x^2-16 is a parabola whose roots are at -4,4 and vertex is at 16 when x=0 x^2+4x+4x+16 is a parabola whose root is at -4, this is also its vertex where y=0 and x=-4 That's about all i can think of...